7. Find the area of the region enclosed by the cardioid r = 2-2 sine. please show how to find the interval!!!!!!!! I repeat it's very important that you show your work on how to find the interval!!!!!!!!!!!!
Answer #1

r . 2 - 2 since max & 2-2xE1) = 2+2 = 4 at 0 = 3/2 1 - 2 - 2 x(i) 2 -2 = 0 at c : 7 oers4 => QS 34 but the we limit need for